OpenStudy (anonymous):

METAL! The probability that Roger wins a tennis tournament is 0.75 (event A), and the probability that Stephan wins a golf tournament is 0.64 (event B). The probability that both Roger and Stephan win their respective tournaments is 0.48. Which statement is true?

OpenStudy (anonymous):

Events A and B are independent because P(A and B) = P(A) P(B). Events A and B are independent because P(A and B) = P(A) + P(B). Events A and B are dependent because P(A and B) = P(A) P(B). Events A and B are dependent because P(A and B) = P(A) + P(B). Events A and B are dependent because P(A and B) P(A) P(B).